OpenStudy (anonymous):

Auditions for a play in New York were held with 100 men and women trying out for the lead singer position. If twice the number of men less three times the number of women is equal to 20, how many men auditioned? 44 40 76 64

hero (hero):

m + w = 100 2m - 3w = 20 m = 100 - w m = (20 + 3w)/2 m = m 2(100 - w) = 20 + 3w 200 - 2w = 20 + 3w 180 = 5w 180/5 = w 36 = w 64 = m

hero (hero):

Nicole, you should check your previous equation to make sure it is correct.